OpenStudy (anonymous):

Solve 3x - 4 ≤ 2 or 2x + 11 ≥ -1.

OpenStudy (anonymous):

3x - 4 equals 1.33333333333 and 2x + 11 = 5.5

OpenStudy (anonymous):

Just taking a stab at it :/

OpenStudy (anonymous):

the answer is between these 3 {x | x ≤ 2} {x | -6 ≤ x ≤ 2} {all reals}

OpenStudy (anonymous):

wait is that an l or a 1

OpenStudy (pinksapphire):

3x-4<=2 or 2x+11>=-1 +4 +4 -11 -11 3x<=6 2x>=-12 3x/3<=6/3 2x/2>=-12/2 x<=2 or x>=-6

OpenStudy (anonymous):

l

OpenStudy (pinksapphire):

I think it's the second one...